On this question, it wants me to drag and drop tiles to match the ordered pairs. The tile choices are:
3x+2y=6
-5x+y=-10
x-4y=8
-6x-5y=30
and the ordered pairs we're matching it to are:
(0,6)
(0,3)
(4,-1)
(1,-5)

Answer by Alan3354(69443)

3x+2y=6
-5x+y=-10
x-4y=8
-6x-5y=30
and the ordered pairs we're matching it to are:
(0,6)
(0,3)
(4,-1)
(1,-5)
===============
Sub for x & y and see if it fits the equation.
eg,
(0,6) and 3x+2y=6
3*0 + 2*6 =? 6
12 =? 6
No, that's not one of them.
==================
Do the others the same way.